# Debiased Sample Selection for Combating Noisy Labels

### v1

@@ 3.1 ITEM: noIse-Tolerant Expert Model
Figure 4: The mapping function $\mathcal{S}^{\beta}(\cdot)$ with varying values of $\beta$. Note that $\mathcal{S}(\cdot)$ degrades into a transverse line when $\beta=1$.
Remark. How to update parameters of the expert layer? All selection criteria depend on the model’s performance, including both small-loss family and fluctuation-based methods. Therefore, leveraging expert layers for selecting without updating parameters easily results in poor selection, even leading to a collapse. To fulfill accurate selection on expert layers, we propose a stochastic updating strategy, which randomly assigns a layer from the set of $\mathbbm{F}=\{f,g^{1},...,g^{m}\}$ as the classifier layer $f^{c}$ and the rest are expert layers. Hence, Eq. (3.1) can be rewritten as
$\small\hat{R}_{\text{clean}}(f^{c})=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\textsf{Criterion}(\mathbbm{F}\backslash f^{c},\bm{x}_{i},{\tilde{y}}_{i})\cdot\mathcal{L}(f^{c}(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]}.$ (5)
Since sufficient training processes, we consider that each layer in ITEM contains remarkable performance on selection and prediction.
3.2 Mixed Sampling
To overcome the data bias, we propose a Mixed Sampling strategy to simultaneously learn from head and tail classes in each training iteration. Specifically, supposing a clean set ${\tilde{D}}_{\text{clean}}$ is selected from the ${\tilde{D}_{N}}$, where ${\tilde{D}}_{\text{clean}}\subseteq{\tilde{D}_{N}}$. By calculating the number of instances from each class in ${\tilde{D}}_{\text{clean}}$ and conducting L1 normalization, a vector of class weights $\bm{v}=[\bm{w}_{1},...,\bm{w}_{K}]$ can be obtained. The weight of class $k$ is written as
$\bm{w}_{k}=\texttt{Normalize}(S_{k})=u_{k}\Big{/}\sum\nolimits_{i=1}^{K}u_{i},$ (6)
where $u_{i}$ denotes the quantity of $i$-th class in ${\tilde{D}}_{\text{clean}}$.
@@ 3.2 Mixed Sampling
Based on the weight vector $\bm{v}$, we further propose a reversed weight vector that assigns a larger weight to tailed classes when sampling. Concretely, we construct a mapping function based on the probability density function (PDF) of a Beta distribution. Normally, the Beta distribution’s PDF is controlled by two parameters $\alpha,\beta$. To satisfy both 1) the input interval of $\bm{w}_{k}$ belongs to $\{0,1\}$ and 2) the function contains the monotonically decreasing property, we thus fix $\alpha=1$ and adjust $\beta$ to achieve different slopes. Given a weight $\bm{w}$, the reversed weight ${\bm{\tilde{w}}}$ can be obtained by our mapping function, which is written as
${\bm{\tilde{w}}}=\mathcal{S}^{\beta}(\bm{w})=\frac{1}{\text{B}(1,\beta)}\,(1-\bm{w})^{\beta-1},$ (7)
@@
Algorithm 1 ITEM
Input: Noisy training set $\tilde{D}_{N}$, training epoch $T$, warmup iterations $T_{w}$, a selection criterion $\textsf{Criterion}(\cdot)$, a classifier network with $m$ experts, the mapping function $\mathcal{S^{\beta}(\cdot)}$ with a hyper-parameter $\beta$.
2:  for $t$ = 1, …, $T_{w}$ do
3:     Randomly sample $f^{c}$ from $\mathbbm{F}$, and WarmUp $f^{c}$ on $D_{N}$.
4:  end for
5:  for ${t}$ = 1, …, $T$ do
6:     Select the clean set ${\tilde{D}}_{\mathrm{clean}}$ from $D_{N}$ based on $\textsf{Criterion}(\mathbbm{F})$.
7:     Calculate weighted vector $\bm{v}=[\bm{w}_{1},...,\bm{w}_{K}]$ on ${\tilde{D}}_{\mathrm{clean}}$.
8:     Calculate reversed weighted vector $\bm{\tilde{v}}=[\bm{\tilde{w}}_{1},...,\bm{\tilde{w}}_{K}]$ with Eq. (7).
9:     while $i$ = 1, …, iterations do
10:        Randomly sample a mini batch $B_{{\bm{v}}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\mathrm{clean}}$ according to $\bm{v}$.
11:        Randomly sample a mini-batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x^{\prime}}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\mathrm{clean}}$ according to $\bm{\tilde{v}}$.
12:        Randomly sample $f^{c}$ from $\mathbbm{F}$.
13:        Update the network parameters by minimizing $\frac{1}{b}\sum\nolimits_{i=1}^{b}\mathcal{L}\big{(}f^{c}\big{(}\texttt{MixUp}(\bm{x}_{i},\bm{x^{\prime}}_{i})\big{)},\,\texttt{MixUp}({\tilde{y}}_{i},{\tilde{y}}^{\prime}_{i})\big{)}$.
14:     end while
15:  end for
Training on the tail-focused training batch directly would result in the sparse representation of head classes [52]. To avoid this problem, we separately conduct two times weighted sampling (according to the weighted and reversed weighted vector ${\bm{v,\tilde{v}}}$) in data loader phase. Then, we have two training mini-batches $B_{\bm{v}}$ and $B_{{\bm{\tilde{v}}}}$, which focus on head and tail classes, respectively. By leveraging MixUp strategy [22], the model can learn a more representative feature extractor on the mixed data $\texttt{MixUp}(B_{\bm{v}},B_{{\bm{\tilde{v}}}})$.
Table 1: Test accuracy (mean±$\pm$±std) of methods using ResNet-18/34 on CIFAR-10/100. Note that $\dagger$,$\ddagger$ , ‡ and $\sharp$ denote three selection criteria, i.e., small-loss detection with loss threshold [10], small-loss detection with Gaussian Mixture Model [12], and fluctuation-based noise filtering [11]. Bold values denote the best the second best performance.
@@
CIFAR-10 Cross-Entropy [53] 85.00 $\pm$ 0.43% 79.59 $\pm$ 1.31% 85.92 $\pm$ 1.09% 79.91 $\pm$ 1.41% 82.61
MentorNet [54] 80.49 $\pm$ 0.11% 77.48 $\pm$ 3.45% 79.12 $\pm$ 0.42% 70.27 $\pm$ 1.52% 76.84
@@
CIFAR-100 Cross-Entropy [53] 57.59 $\pm$ 2.55% 45.74 $\pm$ 2.61% 59.85 $\pm$ 1.56% 43.74 $\pm$ 1.54% 51.73
MentorNet [54] 52.11 $\pm$ 0.10% 35.12 $\pm$ 1.13% 51.73 $\pm$ 0.17% 40.90 $\pm$ 0.45% 44.96
@@
3.3 Practical Training
Despite the training objective of selection-based methods being summarized as a joint framework (in Eq. (5)), the training procedure is practically decoupled into two stages that are iteratively proceeding, 1) select clean samples via the preset criterion and 2) train the network on the selected set. Our training manner also obeys this framework.
Specifically, supposing a selection criterion $\textsf{Criterion}(\cdot)$ such as small-loss [10] or fluctuation-based noise filtering [11] and a network with $m$ size expert layer. In each training epoch, the clean set ${\tilde{D}}_{\text{clean}}$ is selected. To calculate the number of each class in ${\tilde{D}}_{\text{clean}}$, we have a weighted vector $\bm{v}$. Then, the reversed vector $\bm{\tilde{v}}$ is obtained via the mapping function (in Eq. (7)). In each training iteration, we respectively sample a head-focused batch $B{{\bm{v}}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ and a tail-focused batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x^{\prime}}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\text{clean}}$ according to ${\bm{\tilde{v}}}$, where $b$ denote the size of the mini-batch. By randomly selecting a layer from $\mathbbm{F}$ as the classifier layer $f^{c}$, the parameter’s update can be achieved by minimizing the following training loss
$L^{\mathrm{train}}=\frac{1}{b}\sum\nolimits_{i=1}^{b}\mathcal{L}\big{(}f^{c}\big{(}\texttt{MixUp}(\bm{x}_{i},\bm{x^{\prime}}_{i})\big{)},\,\texttt{MixUp}({\tilde{y}}_{i},{\tilde{y}}^{\prime}_{i})\big{)},$ (8)
@@ 3.3 Practical Training
4 Experiments
Datasets. We assess the performance of ITEM on two noise-synthetic datasets CIFAR-10 and CIFAR-100 [48], and two real-world noisy datasets CIFAR-10N [57] and Clothing-1M [58]. To manually generate the noise, we follow the work [8] and construct symmetric and instance-dependent label noise on CIFAR-10 and CIFAR-10. Specifically, given a noise ratio $\rho$, symmetric label noise randomly translates the true label to a wrong label with probability $\rho$, while instance-dependent label noise assumes a corresponding relationship exists in input features [59]. In experiments, we set the $\rho\in\{20\%,40\%\}$. For testing real-world performance, we employ CIFAR-10N, a human-annotated dataset with varying noise conditions, including worst and random 1&2&3. Besides, Clothing-1M [58], a large-scale dataset that is collected from the website and contains one million training data with 14 categories, is also utilized.
Compared methods. On CIFAR-10 & 100, we compare ITEM with prevailing methods that can be roughly divided into three parts, including 1) sample selection methods (small-loss), MentorNet [54], JoCoR [8], and Me-Momentum [16], 2) sample selection methods (fluctuation), SFT [11] and Late Stopping [19], 3) others, Cross-Entropy [53], Joint Optim [55], and PES [56]. Note that all reported results in this paper are collected from SFT and Late Stopping. On the real-world dataset CIFAR-10N, we compare ITEM with Co-teaching+ [14], Peer Loss [60], CAL [61], DivideMix [12], ELR+ [15], and CORES* [59]. The results are collected from the literature [57].
Experimental setup. We keep the convention from [11, 19] and adopt ResNet-18 and ResNet-34 for CIFAR-10 and CIFAR-100, respectively. For all noisy conditions on CIFAR-10 & 100, we leverage an SGD optimizer with the momentum 0.9 and the weight-decay 1e-3 to train our network. The total training epoch is set as 200. The initial learning rate is 0.02 and decayed with the factor 10 at the 100-th and 150-th epoch. For real-world noise, ResNet-34 and a pre-trained ResNet-50 are used for Clothing-1M and CIFAR-10N, respectively. More details can be found in the Table in Appendix C.
Hyperparameter setup. Our framework contains two main hyperparameters, i.e., the number of expert $m$ in the network architecture, and the slope parameter $\beta$ in the mapping function. We keep $\beta=2$ for all experiments and set $m=2,4$ for CIFAR and Clothing-1M, respectively.
4.1 Comparison Results
Results on CIFAR-10 and CIFAR-100. Since our framework is not restricted by different selection criteria, we test our proposal on three types of selection criterion, including small-loss with loss threshold [10], small-loss with Gaussian Mixture Model [12] and fluctuation-based noise filtering [11]. Each trial is repeated five times with different seeds. Mean values with standard deviation are recorded.
@@ 4.1 Comparison Results
[…]
Methods CIFAR-10N
Worst R. 1 R. 2 R. 3
CE 77.69 85.02 86.46 85.16
Co-teaching+ 83.26 89.70 89.47 89.54
Peer Loss 82.00 89.06 88.76 88.57
CAL 85.36 90.93 90.75 90.74
Late Stoping 85.27 - - -
$\ddagger$ ITEM 91.15 95.12 95.03 95.17
DivideMix 92.56 95.16 95.23 95.21
ELR+ 91.09 94.43 94.20 94.34
CORES* 91.66 94.45 94.88 94.74
$\ddagger$ ITEM* 93.14 96.08 96.54 96.71
[…]
Methods Test acc. (%) Methods Test acc. (%)
CE 64.54 Co-teaching 69.21
Forward 69.84 SFT+ 75.08
JoCoR 70.30 DMI 72.46
Joint Optim 72.23 PES (semi) 74.85
DivideMix 74.76 ELR+ 74.81
$\ddagger$ ITEM 74.69 $\ddagger$ ITEM* 75.02
Results on CIFAR-10N. We evaluate the effectiveness of ITEM on a human-annotated noisy dataset CIFAR-10N which contains four noisy conditions, i.e., Worst and Random 1&2&3 [57]. Besides, following previous selection-based approaches [12], we discard labels of unreliable samples and train the model with them by semi-supervised learning (SSL) for greater generalization, which is represented as ITEM*. Appendix A shows the detailed algorithm flowchart for ITEM*.
The results can be found in Table 2. First, our method achieves remarkable performance on both two settings (w/ or w/o SSL). When training without SSL, ITEM outperforms previous methods by a large margin. Compared with Late Stopping, a newly proposed method, ITEM achieves 5.88% improvements on worst-labels. When training with SSL, ITEM is also superior to the previous SOTA method DivideMix. It is worth noting that even training without SSL, ITEM surprisingly gains greater performance than those methods that leverage SSL approaches.
Results on Clothing-1M. Going beyond the small-scale benchmark such as CIFAR-10 & 100, we also test the performance of ITEM and ITEM* on a large-scale dataset Clothing-1M, which contains roughly 38.49% noisy labels. The results are shown in Table 3. We can observe that both ITEM and ITEM* outperform existing methods by a meaningful margin. For example, ITEM* achieves 0.17% improvements compared with the second-best method PES. The result verifies that our method is effective for enhancing model generalization in large-scale real-world scenarios.
[…]
$\xi$ 10
$\rho$ 10% 20% 30% 40% 50%
CIFAR-10 Co-teaching 80.30 78.54 68.71 57.10 46.77
Sel-CL+ 86.47 85.11 84.41 80.35 77.27
RCAL 88.09 86.46 84.58 83.43 80.80
Ours 87.32 86.31 84.91 84.72 81.04
CIFAR-100 Co-teaching 45.61 41.33 36.14 32.08 25.33
Sel-CL+ 55.68 53.52 50.92 47.57 44.86
RCAL 57.50 54.85 51.66 48.91 44.36
Ours 56.10 55.93 53.07 48.29 45.17
Table 5: Ablation study of each component in $\ddagger$ ITEM with varying noise conditions.
Methods CIFAR-10 CIFAR-100 CIFAR-10N
Sym 40% Inst. 40% Sym 40% Inst. 40% Worst
w/o robust multi-expert network (training on normal ResNet) 90.75 91.44 73.01 70.04 89.72
w/o mixed sampling (training on a random sampled batch) 88.77 89.91 71.09 68.54 87.66
w/o mixup (directly training on $B{{\bm{v}}}$ and $B{{\bm{\tilde{v}}}}$) 90.19 90.35 72.50 69.11 89.10
Ours 92.81 93.13 74.90 71.44 91.15
(a) The number of expert layers $m$.
(b) The slope parameter $\beta$.
Figure 5: Sensitivity analyses for hyperparameters selection.

### v3

@@ 3.2 Beta-Resampling
Figure 4: Mapping function $\mathcal{S}^{\beta}(\cdot)$ with different values of $\beta$.
To overcome the data bias, we propose a mixed beta-sampling strategy that promotes the sampling frequency for tail classes in the selected clean set when sampling. Thus, the model can learn from more knowledge from the tailed classes, contributing to mitigating the data bias.
Specifically, suppose a clean set ${D}_{\text{clean}}$ is selected from the ${D_{N}}$, where ${D}_{\text{clean}}\subseteq{D_{N}}$. The set of samples labeled as class $k$ in ${D}_{\text{clean}}$ is ${D}_{\text{clean}}^{k}$. By introducing L1 normalization, a vector of class frequencies $\bm{v}$ can be obtained,
$\bm{v}=[w_{1},w_{2},...w_{K}],\,\text{where}\,\,w_{k}=\frac{\|{D}_{\text{clean}}^{k}\|}{\|{D}_{\text{clean}}\|}.$ (6)
Here, we can regard the class frequency as the sampling weight, i.e., the class with a larger frequency will be sampled with a greater weight during the data loading phase.
@@
To mitigate the data bias, which leads the model to learn from the majority classes mainly, we propose a mapping function that maps the lower frequency to a higher weight when data sampling. Concretely, this mapping function stems from the probability density function (PDF) of a Beta distribution. Normally, the Beta distribution’s PDF is controlled by two parameters $\alpha,\beta$. To satisfy both 1) the input interval of ${w}_{k}$ belongs to $[0,1]$ and 2) the function contains the monotonically decreasing property, we thus fix $\alpha=1$ and adjust $\beta$ for different slopes. Given a weight $w_{i}$, the reversed weight ${\tilde{w}}_{i}$ is written as
${\tilde{w}_{i}}=\mathcal{S}^{\beta}(w_{i})=\frac{1}{\text{B}(1,\beta)}\,(1-w_{i})^{\beta-1},$ (7)
@@
The mapping function $\mathcal{S}^{\beta}(\cdot)$ for varying values of $\beta$ is illustrated in Figure 4. It can be observed that as the input weight increases, the output value decreases. Consequently, this mapping function enables the transformation of class frequencies in Eq. 6 into a tail-focused weighted vector ${\bm{\tilde{v}}}=[\tilde{w}_{1},\ldots,\tilde{w}_{K}]$. By weighted sampling with ${\bm{\tilde{v}}}$, a tail-classes focused training batch $B_{{\bm{\tilde{v}}}}$ can be obtained. Training the model on $B_{{\bm{\tilde{v}}}}$ can effectively mitigate the data bias.
Table 1: Test accuracy (mean±$\pm$±std) of methods using ResNet-18/34 on CIFAR-10/100. Note that $\dagger$,$\ddagger$ , ‡ and $\sharp$ denote three selection criteria, i.e., small-loss selection with loss threshold [11], small-loss selection with Gaussian Mixture Model [21], and fluctuation-based noise filtering [39]. Bold values denote the best the second best performance.
@@
CIFAR-10 Cross-Entropy 85.00 $\pm$ 0.43% 79.59 $\pm$ 1.31% 85.92 $\pm$ 1.09% 79.91 $\pm$ 1.41% 82.61
@@
CIFAR-100 Cross-Entropy 57.59 $\pm$ 2.55% 45.74 $\pm$ 2.61% 59.85 $\pm$ 1.56% 43.74 $\pm$ 1.54% 51.73
@@
3.3 Stochastic Classifier Training
Since the noise filter and the optimizer are disentangled in Eq. (3.1), largely mitigating the training bias, the parameter update of expert layers is unreachable. Considering almost all existing selection criteria rely on the model’s performance, dynamically updating the parameters of expert layers is necessary. To fulfill this goal, we propose a stochastic training strategy, which randomly assigns a layer from the set of $\mathbbm{F}=\{f,g^{1},...,g^{m}\}$ as the classifier layer $f^{c}$ and the rest are expert layers. Hence, Eq. (3.1) can be rewritten as
$\hat{R}_{\text{clean}}(f^{c})=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\mathrm{Criterion}$ $(\mathbbm{F}\backslash f^{c},\bm{x}_{i},{\tilde{y}}_{i})\cdot$
$\mathcal{L}(f^{c}(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]}.$ (8)
Since the aim-to-optimized layer $f^{c}$ is initialized from the set $\mathbbm{F}$ in each iteration, the performance of each layer can be ensured under sufficient training processes.
Despite a highly integrated framework in Eq. (3.3) proposed, the training manner is practically decoupled into three stages that are proceeding iteratively, 1) select clean samples via the preset criterion based on the ensemble result of all experts, 2) randomly select an optimization layer, and 3) train this layer on the selected set.
Considering training on the tail-focused training batch directly would result in the sparse representation of head classes [55], we separately conduct two times weighted samples (according to the weighted and reversed weighted vector ${\bm{v,\tilde{v}}}$) in each data-loading phase, which focus on head and tail classes, respectively. By leveraging the MixUp strategy [53], the model can learn a more representative feature extractor on the mixed data. Specifically, in each training iteration, we respectively sample a head-focused batch $B_{\bm{v}}=\{(\bm{x}_{i},{\tilde{y}}_{i})\}_{i=1}^{b}$ and a tail-focused batch $B_{{\bm{\tilde{v}}}}=\{(\bm{x^{\prime}}_{i},{\tilde{y}^{\prime}}_{i})\}_{i=1}^{b}$ from ${\tilde{D}}_{\text{clean}}$ according to ${\bm{\tilde{v}}}$, where $b$ denotes the size of the mini-batch. By randomly selecting a layer from $\mathbbm{F}$ as the optimized-classifier layer $f^{c}$, the parameter’s update can be achieved by minimizing
$L^{\mathrm{train}}=\frac{1}{b}\sum\nolimits_{i=1}^{b}\mathcal{L}\big{(}f^{c}\big{(}\mathrm{MixUp}(\bm{x}_{i},\bm{x^{\prime}}_{i})\big{)},\,\mathrm{MixUp}({\tilde{y}}_{i},{\tilde{y}}^{\prime}_{i})\big{)},$ (9)
where $\mathrm{MixUp}(,)$ is represented as $\mathrm{MixUp}(a,b)=\gamma\cdot a+(1-\gamma)\cdot b$ and $\gamma$ is a trade-off coefficient randomly sampled from a beta distribution $\texttt{beta}(\delta,\delta)$. The algorithm flowchart of ITEM is shown in Algorithm 1 (see Appdx).
4 Experiments
Datasets. We assess the performance of ITEM on two noise-synthetic datasets CIFAR-10 and CIFAR-100 [18], two human-annotated datasets CIFAR10N and CIFAR100N [38], and three real-world noisy benchmarks including Clothing-1M [46], Food-101N [20], and Webvision [23].
Baselines. On CIFAR-10 and CIFAR-100, we compare ITEM with prevailing methods that can be roughly divided into three parts, including 1) small-loss based selection, JoCoR [37], and Me-Momentum [3], 2) fluctuation-based selection, SFT [39] and Late Stopping [51], 3) others, Cross-Entropy, Joint Optim [34], and PES [4]. Reported results in this paper are collected from SFT and Late Stopping. More information is shown in Appdx.
@@ 4 Experiments
Methods CIFAR10N CIFAR100N
Worst R1 R2 R3
Co-teaching+ [50] 83.26 89.70 89.47 89.54 60.37
Peer Loss [27] 82.00 89.06 88.76 88.57 57.59
CAL [57] 85.36 90.93 90.75 90.74 61.73
Late Stoping [51] 85.27 - - - -
$\ddagger$ ITEM 91.15 95.12 95.03 95.17 69.47
DivideMix [21] 92.56 95.16 95.23 95.21 71.13
ELR+ [26] 91.09 94.43 94.20 94.34 66.72
CORES* [8] 91.66 94.45 94.88 94.74 61.15
DPC [58] 93.82 95.97 95.92 95.90 71.42
$\ddagger$ ITEM* 93.14 96.08 96.54 96.71 72.40
Implementation details. Our code implements utilize Pytorch 1.9.0 and all experiments are run on a single RTX 4090 GPU. We keep the convention from [39, 51] and adopt ResNet-18 and ResNet-34 for CIFAR-10 and CIFAR-100, respectively. For all noisy conditions on CIFAR-10 & 100, we leverage an SGD optimizer with the momentum $0.9$ and the weight-decay $1\times 10^{-3}$ to train our network. The total training epoch is set as 200. The initial learning rate is 0.02 and decayed with the factor 10 at the 100-th and 150-th epoch. For real-world noise, we adopt a ResNet-34 for CIFAR-N and a pre-trained ResNet-50 for Clothing-1M and Food101N, and Inception-ResNet-V2 [33] for WebVision. Other training set-ups like the number of training epochs, the learning rate and its adjustment, the optimizer, and so on keep the same as [25].
Hyper-parameters. Our framework contains two main hyper-parameters, i.e., the number of expert $m$ in the network architecture, and the slope parameter $\beta$ in the mapping function. We keep $\beta=2$ for all experiments and set $m=4$ for CIFAR and $m=2$ for three open-world datasets.
4.1 Results on Synthetic Noisy Datasets
We conduct comparison experiments with two synthetic noise types on CIFAR-10 & 100, where we manually construct different noisy types on these two datasets. Since our framework is agnostic to selection criteria, we test our proposal on three types of selection criteria. Experimental results are shown in Table 1.
@@ 4.1 Results on Synthetic Noisy Datasets
Table 3: Test accuracy (%) comparison of previous methods on three real-world noisy benchmarks. Top-1 and Top-5 test accuracy are reported for the dataset WebVision.
Methods Food101N Clothing1M WebVision
Top-1 Top-5
Co-teaching [11] 83.73 67.94 63.58 85.20
JoCoR [37] 84.04 69.06 63.33 85.06
CDR [43] 86.36 66.59 62.84 84.11
ELR+ [26] 85.77 74.81 77.78 91.68
DivideMix [21] 86.73 74.76 77.32 91.64
SFT+ [39] - 75.00 77.27 91.50
CoDis* [45] 86.88 74.92 77.51 91.95
SURE [25] 88.00 75.10 78.94 92.34
$\ddagger$ ITEM* 88.14 75.08 80.20 93.07
4.2 Results on Real-World Noisy Datasets
We test ITEM’s performance on two human-made noisy datasets and three website noisy datasets. Considering that semi-supervised learning can utilize the sample discarded by the sample selection framework to regularize the model’s learning [21], we integrate ITEM with the Pseudo-labeling technique [19] to further improve the performance.
1) Results on CIFAR10N & 100N. The results are shown in Table 2. First, our method achieves remarkable performance on both two settings (w/ or w/o SSL). When training without SSL, ITEM outperforms previous methods by a large margin. Compared with Late Stopping, a newly proposed method, ITEM achieves 5.88% improvements on worst-labels. When training with SSL, ITEM is superior to the previous SOTA method, DivideMix. Even training without SSL, ITEM surprisingly gains greater performance than those methods that leverage SSL approaches.
2) Results on Food101N, Clothing1M, and WebVision. The results are shown in Table 3. On Food101N, our proposal achieved the best performance, outperforming SURE (the method with the second-best performance) [25] by 0.14%. On Clothing1M, the performance of our proposal is not the best but considerably competitive. Compared with the best performance, the 75.1% test accuracy, ITEM obtained a closed score, i.e., 75.08%. On WebVision, we keep the convention from the work [45] that tests the performance on the validation set of WebVision. ITEM significantly improved Top-1 accuracy by nearly 1.3%.
The result verifies that our method effectively enhances model generalization in large-scale real-world scenarios.
Figure 5: Visualization of debias learning in a class-level. We selected a ResNet-18 as the backbone and compared class-level prediction results of various methods on CIFAR-100 with four noise types. “CE (clean)" denotes training the model on the completely clean set (50k samples in total). The index of classes is sorted according to the class-level accuracy.
Figure 6: Robustness comparison of different networks under two selection criteria. Our proposed MoE structure significantly mitigates the adverse effect of noisy labels.